OpenStudy (anonymous):

Jamal noticed that 3/5 of the children on the playground where boys. Which could be the numbers of boys and girls on the playground? A. 3 boys 5 girls. B. 5 boys 3 girls. C. 8 boys 5 girls. D. 15 boys and 10 girls.

OpenStudy (anonymous):

So if it's 3 boys then it's 2 girls because 5-3 = 2

OpenStudy (anonymous):

2/5 are girls can you figure out the answer from that ?

OpenStudy (anonymous):

15+10= 25 15 are boys so 15/25 and then 15/25= 3/5.

OpenStudy (anonymous):

oh sorry I was away using the bathroom lol

OpenStudy (anonymous):

okay so what's the answer ?

OpenStudy (anonymous):

3 boys and 5 girls